OpenStudy (anonymous):

Find a vector parametrization of the circle of radius 6 in the xy-plane, centered at the origin, oriented clockwise so that the point (6,0) corresponds to t = 0 and the point (0,-6) corresponds to t = 1.

OpenStudy (anonymous):

I keeping getting \[(6\sin (\pi t))i+(6\cos(\pi t))j\]